I have a problem I don't have a clue on how to solve? If I could get some help I'd appreciate it very much. My last question was never answered that I can see, so if someone could help me I'd be very greatful.
Solve: |7x-4|=10 ?
Thank you

Answer by mathslover(157)

We have , |a| = a for a > 0
|a| = -a for a < 0
for the expression |7x -4 | = 10
for 7x -4 > 0 or x > 7/4 we have
7x -4 =10
7x = 14.
x=2 which is > 7/4 . hence x=2 is one solution
for 7x -4 < 0 or x < 4/7 we have
-(7x - 4) =10
or -7x =6
x= -6/7 which satisfies x < 4/7
therefore -6/7 is the other solution
Hence solution set for the equation is { 2, -6/7}